OpenStudy (maheshmeghwal9):

What is the mistake in the computation of: - \[\color{blue}{1=\sqrt{1}=\sqrt{(-1)(-1)}=\sqrt{-1} \times \sqrt{-1}=i \times i =i^2=-1.}\]

Parth (parthkohli):

-1 * -1 = 1

Parth (parthkohli):

Look at the fourth step

Parth (parthkohli):

\(\Rightarrow \sqrt{(-1)(-1)} = \sqrt{1} \)

Parth (parthkohli):

No need to write \(\sqrt{-1} \times \sqrt{-1}\)

Parth (parthkohli):

When you write something in this form, it all falls apart.

OpenStudy (maheshmeghwal9):

it is nt the answer @ParthKohli :/

Parth (parthkohli):

\(\sqrt{(-1) \times (-1)} = \sqrt{1}\)

OpenStudy (anonymous):

No, Parth, it is like this: you can only write sqrt(a)*sqrt(b)=sqrt(ab) if and ONLY if a & b are not negative at the same time.

Parth (parthkohli):

\(\sqrt{a} \times \sqrt{b} = \sqrt{ab}\) Let's try it when both a and b are negative. \(\sqrt{-25} \times \sqrt{-4}\) \(\Rightarrow \sqrt{(-25)(-4)} \) \(\Rightarrow \sqrt{100} \) \(\Rightarrow10\) But this is wrong

Parth (parthkohli):

Yes, rajath

OpenStudy (anonymous):

I sure wish I knew Latex :'(

Parth (parthkohli):

There's another way to do this: \(\sqrt{-25} = 5i\) \(\sqrt{-4} = 2i\) \(5i \times 2i = 10i\)

Parth (parthkohli):

\(10i^2\)*

Parth (parthkohli):

\(10 \times -1\) \(\Rightarrow -10\)

OpenStudy (maheshmeghwal9):

hey hey going so long i want only ur conclusion about the mistake in the question.

Parth (parthkohli):

The fourth step is wrong as \(\sqrt{a} \times \sqrt{b} \ne \sqrt{ab}:a,b=-ve\)

OpenStudy (maheshmeghwal9):

so it would be \[\sqrt{a}\times \sqrt{b}=-\sqrt{ab} \space ; \space a,b=-ve.\] so can u give me the whole & correct solution again plz:)

Parth (parthkohli):

Okay, so we were at this step: \(\sqrt{(-1)(-1)}\) But wait, we can simplify the radical. -1 * -1 = 1 \(\sqrt{-1 \times -1} = \sqrt{1} = 1\)

OpenStudy (anonymous):

no, parth.

Parth (parthkohli):

Yes, Rajat.

OpenStudy (maheshmeghwal9):

bt it is not \[-\sqrt{ab}\] @ParthKohli

Parth (parthkohli):

Actually it is plus or minus 1, but minus one is an extraneous solution.

Parth (parthkohli):

What do you mean by \(-\sqrt{ab}\)?

OpenStudy (maheshmeghwal9):

it is the correct solution of this question:)

OpenStudy (anonymous):

take sqrt(-1)=i & then proceed. So, sqrt(-6) for example will be sqrt(6)*i

Parth (parthkohli):

Yeah, of course. But when you could simplify the thing inside the radical, then why not?

OpenStudy (anonymous):

No, you are going against your own answer.

Parth (parthkohli):

-1 * -1 = 1 right? so \(\sqrt{(-1)(-1)}= \sqrt{1}\)

OpenStudy (anonymous):

But not inside a radical.

Parth (parthkohli):

Can you be more specific, Rajath?

OpenStudy (maheshmeghwal9):

\[\color{blue}{1=\sqrt{1}=\sqrt{(-1)(-1)}=\sqrt{-1} \times \sqrt{-1}=i \times i \sqrt{1} =i^2=-1.}\]well then also i m getting this @rajathsbhat

OpenStudy (anonymous):

Just refer to my first response, Parth.

Parth (parthkohli):

I know, that is what I am saying.

OpenStudy (anonymous):

You can't split the '1' INSIDE the radical as (-1)(-1).

Parth (parthkohli):

You need not stretch a problem like \(\sqrt{60.5 \times 2}\) to \(\sqrt{60.5} \times \sqrt{2}\), as that may cause problems.

Parth (parthkohli):

You could compute 60.5 * 2 which is 121 so \(\sqrt{121} = 11\)

OpenStudy (maheshmeghwal9):

well it is an IIT question. bt u r nt giving me satisfactory answer @ParthKohli @rajathsbhat :/ Plz give me stepwise solution i wanna see wt do u wanna say actually.

OpenStudy (anonymous):

The point is, if you have a +ve numer inside the radical, it can only have come from multiplying 2 +VE numbers.

Parth (parthkohli):

Can you please be more specific as to what we can do, Mahesh?

Parth (parthkohli):

We can split it into (-1)(-1), but we can't make it go like \(\sqrt{-1} \times \sqrt{-1}\)

OpenStudy (maheshmeghwal9):

ya just as i did the whole solution. \[\color{blue}{1=\sqrt{1}=\sqrt{(-1)(-1)}=\sqrt{-1} \times \sqrt{-1}=i \times i \sqrt{1} =i^2=-1.}\]i want the correct one from u @ParthKohli @rajathsbhat

Parth (parthkohli):

Okay, I explained it already. \(\sqrt{a} \times \sqrt{b} \ne \sqrt{ab}:a,b\in -ve\)

OpenStudy (maheshmeghwal9):

ok

Parth (parthkohli):

Do you need any more explanation of this?

OpenStudy (anonymous):

I'm telling you this exact "paradox" was raised by my lecturer when I was undergoing IIT JEE coaching. The answer to this problem IS MY FIRST RESPONSE.

OpenStudy (maheshmeghwal9):

wait a minute i have the solution at the end of my book i m just typing that.

Parth (parthkohli):

No need to gobble this all up from your book. Do not believe in these fake IIT books.

OpenStudy (anonymous):

You see, Parth, your 3rd latest response just doesn't allow for you to write sqrt(−1)(−1)=sqrt1

OpenStudy (maheshmeghwal9):

lol parth The answer is: - The formula \[\sqrt{a}\sqrt{b}=\sqrt{ab}\]is applicable only when at least one from a & b is positive. otherwise \[\sqrt{-a}\sqrt{-b}=-\sqrt{ab} \space ; \space a,b>0\]Well i want answer acccording to this @ParthKohli & @rajathsbhat

OpenStudy (anonymous):

This is exactly what I've been trying to sat from the beginning.(My first response=just worded differently). D: I think Parth was saying the same thing too....

OpenStudy (anonymous):

say* not sat.

OpenStudy (maheshmeghwal9):

bt actually parth was asking \[-\sqrt{ab}=?\]well that is the answer can u give me now the correct method of the question.?

OpenStudy (anonymous):

can you explain that again? I don't understand your ques.

Parth (parthkohli):

@maheshmeghwal9 Isn't that what I said?

OpenStudy (maheshmeghwal9):

i wanna say that are u seeing the process of sqrt{1} in our original question can u do it now correctly ? & if yes plz give me full solution:)

Parth (parthkohli):

There is not just one notation in this world

OpenStudy (maheshmeghwal9):

actually correct the mistake & give the question in correct form again omg

OpenStudy (anonymous):

The original question is flawed in going from step 2 to 3.

Parth (parthkohli):

No actually step 2 -> 3 is still acceptable, but then 3 -> 4 was applied wrong.

OpenStudy (maheshmeghwal9):

? getting more confused aarrghhhhh. @ujjwal Plz help:)

OpenStudy (anonymous):

NO.NO.NO....

OpenStudy (anonymous):

1=sqrt(1) That's all. You can't do any thing furthur. If you want you can only write sqrt(1)=sqrt[(1)(1)]=sqrt[(1)(1)(1).....] That's all & that's it.

OpenStudy (maheshmeghwal9):

so i gt u now clearly @rajathsbhat but the main problem is that why can't we do it by taking -1 x -1 =1.

OpenStudy (anonymous):

Let me write something in paper & post it> I don't know Latex & I'm finding it hard to explain my thoughts.

OpenStudy (maheshmeghwal9):

k find my mistake now it is now improved version lol @rajathsbhat \[\color{blue}{1=\sqrt{1}=\sqrt{(-1)(-1)}=\sqrt{-1} \times \sqrt{-1}=i \times i \sqrt{1 \times 1} =i^2\sqrt{1}=-1.}\]

OpenStudy (anonymous):

wait for 5.

OpenStudy (maheshmeghwal9):

ok:)

OpenStudy (anonymous):

\[1=\sqrt{1}\neq \sqrt{(-1)(-1)}\]

OpenStudy (anonymous):

do you get it?

OpenStudy (ujjwal):

At least i don't get it. @rajathsbhat

OpenStudy (maheshmeghwal9):

is there any reason why it is nt like this?

OpenStudy (maheshmeghwal9):

@rajathsbhat

OpenStudy (anonymous):

I'm sorry terrible mistake.\[1=\sqrt{1}= \sqrt{(-1)(-1)}\neq \sqrt{-1}\sqrt{-1}\]

OpenStudy (ujjwal):

Is that just a postulate which can't be questioned or it has some logical reasoning?

OpenStudy (maheshmeghwal9):

then it is like this \[1=\sqrt{1}= \sqrt{(-1)(-1)}= \sqrt{1}i\sqrt{1}i=1i^2=-1\]

OpenStudy (anonymous):

because the opposite is not true. ie..\[\sqrt{-1}\sqrt{-1}\neq \sqrt{(-1)(-1)}\]

OpenStudy (maheshmeghwal9):

no; it is a sensible question & correct too from IIT @ujjwal

OpenStudy (anonymous):

Now that I can use Latex, I'll explain one big thing that summarizes everything I said. Is that OK? Sorry for being the dumb-retricethat I am on Latex --.--

OpenStudy (anonymous):

retrice not retrice.

OpenStudy (anonymous):

retrice

OpenStudy (anonymous):

spell-check problem. Trying to type retrice

OpenStudy (maheshmeghwal9):

its ok:)

OpenStudy (ujjwal):

And you are typing retrice, which is not a big issue here. just waiting for your latexed answer.. P.S. 'latexed' is also a spelling mistake... :P .. Infact there is no such word. :P :D

OpenStudy (anonymous):

Now, \[\sqrt{a}\sqrt{b}= \sqrt{(a)(b)} \: \: \: \: iff\; \; a\; and\; b\; are \; not \; both\; negative.\]

OpenStudy (anonymous):

Therefore, the reverse is also true.

OpenStudy (anonymous):

If a& b are both negative, \[\sqrt{(a)(b)}\neq \sqrt{a}\sqrt{b}\]

OpenStudy (anonymous):

\[\sqrt{ab}=\sqrt{(a)(b)}= \sqrt{ab}\] if both a & b are -ve. There is nothing more you can do about it.

OpenStudy (ujjwal):

\[{\color{DarkOrange} {Thanks -Rajathsbhat}}\]

OpenStudy (anonymous):

Sure thing. :)